OpenStudy (anonymous):

find the difference of 3/5 - 3/8

jhonyy9 (jhonyy9):

3/5 -3/8 =(8*3-3*5)/40 =(24-15)/40=9/40

OpenStudy (anonymous):

find the LCD, which is 40 giving you 24/40-15/40 =9/40

jhonyy9 (jhonyy9):

because 40 is the denominator common for 8 and 5